OpenStudy (anonymous):

Which of the following equations would graph a line parallel to 3x − y = 7?  (1 point)

OpenStudy (anonymous):

y = 7x − 3 y = −3x + 7 y = −7x + 3 y = 3x

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

No. What is the slope of the line 3x - y = 7. Show me your work.

OpenStudy (anonymous):

Hint: WQrite the line in y = mx + b form, so m is the slope.

OpenStudy (anonymous):

Give up?

OpenStudy (owlcoffee):

So let's see something about the Line equiation. We Know it has this form: \[y=mx+b\] where m is the slope and b is the OY coordinate. By condition of parallelism, we know that the slopes must be the same! So, what do we require for a line to be parallel to an other?

OpenStudy (anonymous):

y=3x+7?

OpenStudy (anonymous):

so its B..?

OpenStudy (owlcoffee):

yes! :)

OpenStudy (anonymous):

Excuse me, that is INCORRECT. It is NOT B.

OpenStudy (anonymous):

Our line has a slope of 3...so we want to pick from the choices, a line whose slope is 3.

OpenStudy (owlcoffee):

Oh my, I lost that -. so yeah, the answe is y=3x

OpenStudy (anonymous):

Correct, y = 3x has a slope of 3!!!!